OpenStudy (anonymous):

find solution for cos2x+cosx=0

OpenStudy (anonymous):

\[\cos^2x+\cos x=0~~?\] or \[\cos2x+\cos x=0~~?\]

OpenStudy (anonymous):

second one

OpenStudy (vijay):

2cos^2(x) + cos x -1 = 0

OpenStudy (vijay):

which is becomes, 2 cos^2 x + 2cos x - cos x -1 = 0

OpenStudy (anonymous):

btw equation is cos2x+cosx=0

OpenStudy (anonymous):

not ^2

OpenStudy (vijay):

yes, cos 2x = 2cos^2(x) -1

OpenStudy (anonymous):

ok yeah

OpenStudy (vijay):

x = pi/4 & 1 more ans

OpenStudy (anonymous):

@vijay what bout cos2x+sinx=0

OpenStudy (vijay):

in this case, think other way cos 2x = 1 - 2sin^2(x)

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

acutally i got it thanks